### v2

# BitDelta: Your Fine-Tune May Only Be Worth One Bit

Abstract
Large Language Models (LLMs) are typically trained in two phases: pre-training on large internet-scale datasets, and fine-tuning for downstream tasks. Given the higher computational demand of pre-training, it’s intuitive to assume that fine-tuning adds less new information to the model, and is thus more compressible. We explore this assumption by decomposing the weights of fine-tuned models into their pre-trained components and an additional delta. We introduce a simple method, BitDelta, which successfully quantizes this delta down to 1 bit without compromising performance. This interesting finding not only highlights the potential redundancy of information added during fine-tuning, but also has significant implications for the multi-tenant serving and multi-tenant storage of fine-tuned models. By enabling the use of a single high-precision base model accompanied by multiple 1-bit deltas, BitDelta dramatically reduces GPU memory requirements by more than 10$\times$, which can also be translated to enhanced generation latency in multi-tenant settings. We validate BitDelta through experiments across Llama-2 and Mistral model families, and on models up to 70B parameters, showcasing minimal performance degradation over all tested settings. Code is avaliable at https://github.com/FasterDecoding/BitDelta.
Machine Learning, ICML
Machine Learning, ICML
1 Introduction
After large-scale pretraining, foundation models are typically fine-tuned for specific downstream tasks (Devlin et al., 2019; Radford et al., 2018, 2019). This pretrain-finetune paradigm has revolutionized machine learning; LLMs have not only proven effective for critical tasks such as instruction following and alignment (Ouyang et al., 2022), but are also performant on a wide array of niche yet highly impactful applications (Xu et al., 2024; Qiu et al., 2023). Through fine-tuning, LLMs are adeptly equipped to align with distinct user preferences or specialized task requirements, showcasing an unprecedented level of adaptability. Thus, the prospect of serving millions of uniquely fine-tuned models, each tailored to individual tasks and user needs, presents a promising vision for the future of machine learning.
Realizing this vision is challenging due to two key reasons: 1) Expensive Storage. Each new fine-tuned model is large, even if we have relatively few base models, making them expensive to store and challenging to manage on disk. 2) Expensive Serving. Distinct fine-tuned models each demand significant GPU memory, making it difficult and expensive to concurrently serve such models without noticeable downtime. To tackle these issues, we decompose the fine-tuned model weights into the weights of the base pre-trained model and a delta induced by the fine-tuning process. By compressing this delta while maintaining model performance, we aim to sidestep the prohibitive costs associated with storage and GPU memory demands.
From the delta decomposition point of view, parameter-efficient fine-tuning (PEFT) methods like LoRA (Hu et al., 2021; Houlsby et al., 2019a; Rebuffi et al., 2017; Dettmers et al., 2023; Chen et al., 2023d) effectively enforce a highly structured and compressed form of delta during fine-tuning, a powerful insight for model serving of PEFT-based fine-tunes. (Sheng et al., 2023) and (Chen et al., 2023b) explore multi-tenant serving of LoRA-based fine-tunes.
Figure 1: Overview of BitDelta. BitDelta applies 1-bit quantization to the weight delta between fine-tuned and base models. For each weight matrix, we quantize its delta as its sign bits and a trainable high-precision scale factor. The scale factor is initialized to achieve the best approximation error in $L_{2}$ norm and further refined with a few distillation steps. BitDelta shows minimal degradation in model performance and reduces memory consumption in multi-tenancy serving by representing multiple fine-tuned models with a single high-precision base model and multiple 1-bit deltas.
Nevertheless, recent work has shown that PEFT methods may not yet match the model quality of full parameter fine-tuning, especially on high resource tasks (Chen et al., 2022), and are fairly sensitive to hyperparameter choice and prompting methods (Niederfahrenhorst et al., 2023). As a result, we notice that among the 2307 LLMs (as of time of writing) on the Open LLM Leaderboard (Beeching et al., 2023) with a valid README file, only $<20\%$ indicate that they exclusively use LoRA. Most models are full parameter fine-tunes, model merges (Yu et al., 2023; Jin et al., 2023; Wortsman et al., 2022) of full parameter fine-tunes, or model merges of LoRA based fine-tunes (which are effectively high-rank).
It is also attractive to approximate general deltas with low-rank matrices post-training. However, experimental results show that this is challenging (Table 1), as deltas from full parameter fine-tunes tend to be fairly high-rank (Figure 2).
Figure 2: CEV plot of a $4096\times 4096$ weight delta between Llama 2-7B and Vicuna-7B v1.5. Deltas from full parameter fine-tuning are fairly high rank, making low-rank approximations difficult.
Instead, we draw from the insight that motivates PEFT methods in general: Given the higher computational demand of pre-training, it’s intuitive to assume that fine-tuning adds less new information to the model, and is thus much more compressible. Indeed, we find that we can efficiently quantize the delta to merely 1 bit with almost no performance drop. We propose BitDelta, an efficient post-training quantization (PTQ) solution that acts on the weight delta between a fine-tuned model and its underlying base model.
BitDelta has two stages: 1) We quantize the delta of each weight matrix into a scaling factor multiplied by a binary matrix. Specifically, we take the sign of the delta to form the binary matrix and initialize the scaling factor as the average of the absolute values of the delta, which minimizes $L_{2}$ norm. 2) We further calibrate the scaling factors through model distillation over a small calibration dataset while keeping the binary matrices frozen. Despite the small number of trainable parameters and training steps, we find that this distillation process is effective in further recovering model quality. BitDelta is extremely efficient compared to other quantization methods; we are able to compress 70B models in roughly 10 minutes, whereas methods like GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2023) may take multiple GPU hours. Our experiments over 17 popular fine-tuned models affirm that BitDelta can be applied across various model types and sizes with minimal impact on performance.
BitDelta opens up opportunities to efficiently serve multiple fine-tuned models with shared servers: By only storing a single full-precision base model, and (dynamically) loading and performing batched inference over multiple 1-bit deltas, we can efficiently represent multiple fine-tuned models. Compared to naively using full precision fine-tuned models, deltas compressed by BitDelta are more than 10$\times$ smaller, consuming less GPU memory, and can therefore be loaded faster. This addresses the storage challenge. Moreover, since LLM inference is memory-bound (Leviathan et al., 2022; Chen et al., 2023a; Cai et al., 2024), the latency of each decoding step is proportional to the GPU memory consumption of the model weights. With an efficient CUDA kernel implementation, we can translate this memory reduction into a latency speedup, similar to other quantization methods (Frantar et al., 2022; Lin et al., 2023). As a preliminary attempt, we implement a 1-bit matrix multiplication kernel in Triton (Tillet et al., 2019) and show that it can improve the multi-tenant serving latency by about 2$\times$, addressing the serving challenge.
Finally, we study a few extensions of BitDelta, where we quantize the base model, and where we iteratively apply BitDelta. Experimental results show that our method is quite general and can be applied to various use cases.
Table 1: Comparison between BitDelta and a SVD based method ($r=16$), with Llama 2-7B and Vicuna-7B v1.5 as the base and fine-tuned models. BitDelta is performant across the board, whereas the SVD-based method fails to sufficiently capture the fine-tuned information.
Model/Method Train Loss TruthfulQA GSM8K MT-Bench Adjusted Average2.1 $\uparrow$
Llama 2-7B 38.96 13.57 60.53
Vicuna-7B v1.5 50.36 19.03 6.04 60.51
BitDelta-Initial 0.41 47.63 19.56 5.67 60.99
BitDelta 0.052 49.97 20.17 5.99 60.68
SVD-Initial 0.78 44.56 15.92 4.45 60.60
SVD 0.34 44.55 14.18 4.98 60.20
2 Related Work
2.1 Full Model Compression
Quantization.
Quantization techniques are widely used to reduce memory consumption and improve LLMs’ generation latency. Xiao et al. (2023) implement a technique that rescales between activations and parameters, effectively mitigating outlier activations to facilitate smoother quantization. Dettmers et al. (2022) develop an approach that decomposes matrix multiplications into 8-bit computations, with an additional 16-bit process for handling outliers. Exploring further, Frantar et al. (2022) introduce a method that iteratively rounds weight columns to 3-4 bits of precision. Similarly, Lin et al. (2023) propose an activation-aware quantization scheme that selectively preserves crucial weights while compressing the majority to 3-4 bits. In a different vein, Kim et al. (2023) devise a sparse, low-precision pattern focusing on a small yet significant set of weights. Lastly, Chee et al. (2023) utilize incoherence processing to quantize model weights to as low as 2 bits with minimal impact on performance.
Pruning.
Pruning also aims to reduce the memory consumption of neural networks. It accomplishes this by pushing certain parameter values to zero, inducing sparsity in the model (LeCun et al., 1989; Han et al., 2015, 2016; Zhu & Gupta, 2017). However, these methods may fail to take advantage of modern hardware like GPUs unless using certain structured sparsity patterns like 2:4 (50%) sparsity (Mishra et al., 2021). Frantar & Alistarh (2023) demonstrate a pruning method on LLMs that successfully utilizes the 2:4 sparsity pattern and achieves a 50% sparsity ratio. It is challenging to obtain higher sparsity while being hardware-friendly.
22footnotetext: Adjusted Average is over ARC, BBH, HellaSwag, WinoGrande, and excludes TruthfulQA, GSM8K, MT-Bench.
2.2 Delta Compression
Parameter-efficient fine-tuning.
Parameter-efficient fine-tuning (PEFT) techniques reduce the number of trainable parameters during fine-tuning, reducing memory and compute demand while achieving promising accuracy. Houlsby et al. (2019b) introduce adapter layers, freezing the base model during fine-tuning. Hu et al. (2021) propose Low-Rank Adaptation (LoRA), which enforces that weight updates during fine-tuning are low rank.
By enforcing structure in the fine-tuning information, PEFT techniques can viewed as compression-aware training. Similar to our method, such compression induced by PEFT techniques can also enhance multi-tenant serving. Chen et al. (2023b) and Sheng et al. (2023) explore this idea by developing scalable multi-tenant systems for LoRA-based fine-tunes, based on customized CUDA kernels that exploit the low-rank structure of the fine-tuned weight deltas.
Early work on post-training delta compression.
Most related to our work, a few studies explore the idea of post-training delta compression by adopting existing compression techniques like GPTQ, unstructured pruning (Han et al., 2016), or even classic lossless compression algorithms. Isik et al. (2023) focus on reducing the delta size to save storage. Yu et al. (2023) utilize pruning to improve model merging applications. Yadav et al. (2023) reduces the size of PEFT modules to save storage. Ryu et al. (2023) combines quantization with low-rank approximation to reduce the delta size. The concurrent and independent work by Yao & Klimovic (2023) also explores using delta compression to improve multi-tenant serving, but focuses more on reducing the model loading time from disk to GPU. Compared to existing work, we offer a much simpler and faster method, BitDelta, achieving a more than 10$\times$ compression ratio while also being friendly to modern accelerators.
3 BitDelta
3.1 Method
BitDelta consists of two stages: 1) We quantize each weight matrix into a scalar multiplied by a binary matrix***In our experiments, we only quantize the linear layers in the Transformer blocks as they contribute most of the parameters and computation.. 2) We further calibrate the scalar factors with distillation. We describe each stage in this section:
1-bit quantization.
Let $W_{\text{base}},W_{\text{fine}}\in\mathbb{R}^{n\times m}$ be weight matrices from the base model and fine-tuned model, respectively. We define the weight delta as $\Delta=W_{\text{fine}}-W_{\text{base}}$, representing the modification in weights post-fine-tuning. For efficient representation, we aim to obtain a binarized estimator of this weight delta, denoted as $\hat{\Delta}$, by encoding its sign bits:
$\hat{\Delta}=\alpha\odot\text{Sign}(\Delta),$ (1)
where
$\text{Sign}(W_{ij})=\begin{cases}+1,&\text{if }W_{ij}>0,\\ -1,&\text{if }W_{ij}\leq 0,\end{cases}$ (2)
and $\alpha$ is a high-precision scaling factor for the entire matrix. To minimize the quantization error in $L_{2}$ norm:
$\left\|\Delta-\hat{\Delta}\right\|_{2}^{2}=\sum_{ij}\left(\left|W_{ij}\right|-\alpha\right)^{2},$ (3)
we take
$\alpha=\frac{1}{nm}\sum_{ij}|W_{ij}|.$ (4)
Surprisingly, we find that the above quantization approach already does quite well and retains most of the fine-tuned models’ performance.
Table 2: BitDelta works on Llama-2 and Mistral families and on a wide range of model sizes ranging from 7B to 70B parameters. BitDelta works for many types of fine-tuned information, including SFT-based methods, RLHF-based methods, and context extension methods (RoPE scaling). Scale distillation is effective; it raises TruthfulQA/GSM8K scores to within 1-2 points of the baseline fine-tune, and MT-Bench scores to within 0.1-0.2 points.
Model Method TruthfulQA GSM8K MT-Bench Adjusted Average2.1 $\uparrow$
Llama 2-7B 38.96 13.57 60.53
Llama 2-7B Chat Baseline 45.32 22.74 6.56 59.81
BitDelta-Initial 41.10 18.27 6.31 60.7
BitDelta 44.95 20.24 6.47 59.88
Vicuna-7B v1.5 16k Baseline 50.38 14.18 6.06 57.50
BitDelta-Initial 45.58 13.95 5.69 58.51
BitDelta 48.75 14.48 6.24 57.64
Llama 2-13B 36.90 22.74 64.68
Llama 2-13B Chat Baseline 43.95 33.13 6.98 63.99
BitDelta-Initial 41.70 33.36 7.06 64.25
BitDelta 43.47 31.92 6.95 63.96
Vicuna-13B v1.5 16k Baseline 50.38 29.72 6.90 57.5
BitDelta-Initial 41.7 26.76 6.60 64.25
BitDelta 48.75 28.73 6.88 57.64
WizardLM-13B v1.2 Baseline 47.17 42.38 6.95 61.61
BitDelta-Initial 44.89 42.08 6.73 61.91
BitDelta 46.67 41.62 6.93 61.86
Llama 2-70B 44.82 52.69 71.81
Llama 2-70B Chat Baseline 52.77 47.61 7.12 68.82
BitDelta-Initial 41.63 42.38 6.85 66.01
BitDelta 51.37 48.82 7.06 69.14
Solar-0-70B Baseline 62.03 56.18 7.07 73.77
BitDelta-Initial 59.08 56.79 6.79 73.14
BitDelta 62.03 56.63 6.82 73.57
Mistral-7B v0.1 42.60 37.76 65.98
Mistral-7B v0.1 Instruct Baseline 55.93 32.75 6.86 60.36
BitDelta-Initial 51.27 38.82 6.54 63.83
BitDelta 55.23 31.54 6.43 61.10
Zephyr-7B-$\beta$ Baseline 55.12 34.34 7.18 65.22
BitDelta-Initial 54.53 40.26 6.70 66.12
BitDelta 58.39 31.92 7.00 66.20
Dolphin 2.2.1 Baseline 54.02 54.28 7.36 67.31
BitDelta-Initial 48.14 50.27 7.10 67.58
BitDelta 54.91 52.84 7.20 66.97
Table 3: Comparison of model responses from Zephyr-7B-$\beta$ for Question 9 in MT-Bench, a concise advertisement task. BitDelta-Initial is unable to follow the instructions, producing an advertisement that is overly formal and makes no attempt to adhere to the word limit. With the addition of scale distillation, BitDelta successfully produces a concise, catchy advertisement slightly over the word limit. *Prompt slightly modified for clarity.
Prompt
Please create a slogan about bioenergy that satisfies the following: 1. Make the tone sound casual 2. Embed an advertisement for a company called ”FlexPower” 3. Fewer than 10 words.
BitDelta
”FlexPower’s groundbreaking bioenergy solution: Produce fuel without compromising food production!” (10 words)
This headline is catchy, casual, and highlights the unique selling point of the…
GPT-4 Score: 8
BitDelta-Initial
”FlexPower’s groundbreaking technology unlocks the key to harnessing the power of renewable bioenergy while minimizing ethical concerns”
GPT-4 Score: 4
Scale distillation.
Intuitively, the scaling factor $\alpha$ plays a more significant role in the low-bit regime. Additionally, per-matrix $L_{2}$ weight error is not a perfect measure of degradation in overall model quality. Thus, we further optimize these scales by performing model distillation to align the output logits of the quantized model to that of the original fine-tuned model. More concretely, we freeze the model weights and optimize for the following objective:
$\bm{\alpha}^{*}=\arg\min_{\bm{\alpha}}\mathbb{E}_{x\sim\mathbf{X}}\left[\left\|\mathbf{Z}_{\text{fine}}(x)-\mathbf{Z}_{\text{bin}}(x;\bm{\alpha})\right\|^{2}\right]$ (5)
where $\mathbf{X}$ is a calibration dataset, and $\mathbf{Z}(\cdot)$ are the logits of the respective models. We find that scale distillation is fairly insensitive to choice $\mathbf{X}$, as 1) the process is extremely parameter efficient, and 2) the crucial aspect of the process is to logit match with the fine-tuned model, regardless of the actual text content.
For our experiments, we distill on the C4 dataset (Raffel et al., 2023), which is widely used for pre-training, using 800 samples of length 128, and use the Adam optimizer (Kingma & Ba, 2017) with $lr=10^{-4}$, $\beta=(0.9,0.999)$, $\epsilon=10^{-8}$. 1x80 GB A100 GPU is used to distill 7B and 13B models, and 6x80GB A100 GPUs are used to distill 70B models (2x for finetune, 4x for binarized). Sale distillation is fast; we can compress 70B models in roughly 10 minutes, whereas methods like GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2023) take multiple hours.
3.2 Implication
The ability to compress the delta to merely 1-bit opens up multiple opportunities for improving efficiency, enabling more effective model storage (Isik et al., 2023) – where a single base model can be maintained alongside multiple compressed deltas – and facilitating model hot-swapping  (Chen et al., 2023b; Sheng et al., 2023). With hot-swapping, the base model remains in GPU memory, and compressed deltas are dynamically loaded in accordance to incoming requests. In both cases, the compression ratio can be directly translated into reductions in storage needs and loading times.
Moreover, BitDelta enables the possibility of a multi-tenant serving system like Punica (Chen et al., 2023b) or S-LoRA (Sheng et al., 2023) but for general fine-tuned models instead of just LoRA models. Concretely, we consider the scenario where multiple models fine-tuned from the same base model are served with the same server. This setting greatly exploits the GPU resource and saves each fine-tuned model’s inference cost when their traffic is low or unbalanced. With BitDelta, we can keep one high-precision base model with multiple compressed deltas in the GPU memory. Compared to directly serving multiple fine-tuned models, this approach greatly saves memory consumption.
Since LLM inference follows the memory-bound computation pattern where the generation latency is proportional to the GPU memory used by the model weights, the lower memory consumption also suggests the opportunity to improve the serving latency. For example, Punica and S-LoRA exploit LoRA’s structure and memory saving by developing a CUDA kernel that can efficiently calculate the batched delta-activation product for low rank deltas. Similarly, we decompose the forward pass of each linear layer as follows:
$X^{\prime}_{i}=W_{\text{fine},i}X_{i}\approx W_{\text{base}}X_{i}+\underbrace{\hat{\Delta}_{i}X_{i}}_{\text{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\text{Kernel}}}}$ (6)
where $X_{i}$ and $X_{i}^{\prime}$ represent input and output features to the $i$-th fine-tuned model, and the base model weight and the delta are computed separately. For a batch of requests, $W_{\text{base}}X_{i}$ can be computed with the classic batched GEMM kernel. We implement a fused binary GEMM kernel in Triton (Tillet et al., 2019) that allows us to calculate $\hat{\Delta}_{i}X$ in a batched setting while keeping the 1-bit deltas quantized until they are transferred to the GPU cache. This kernel fuses the dequantization operation with the GEMM calculation, reducing the data moving overhead by a large factor.
Table 4: We apply BitDelta to Llama 2-7B Chat, and find it holds up when the underlying base model is quantized at various levels. FP16 + $\Delta$ outperforms baseline GPTQ across the board, implying that in terms of model quality, we would rather store a single high-precision base model with many 1-bit deltas than store many quantized fine-tuned models. GPTQ + $\Delta$ with Llama 2-7B Chat as the base model also outperforms baseline GPTQ on many evaluations, because the delta diffuses 16-bit information through high precision scaling factors.
Base Model Method TruthfulQA GSM8K MT-Bench Adjusted Average2.1 $\uparrow$
Baseline FP16 45.32 22.74 6.56 59.81
INT8 RTN 45.02 22.29 6.28 59.63
GPTQ 44.92 19.48 5.90 58.67
Llama 2-7B FP16 + $\Delta$ 44.95 20.24 6.47 59.88
INT8 RTN + $\Delta$ 44.71 19.86 6.16 59.85
GPTQ + $\Delta$ 42.52 19.94 6.02 59.22
Llama 2-7B Chat GPTQ + $\Delta$ 44.63 22.14 6.11 59.17
4 Experiments
4.1 Setup
Baselines.
Our primary baselines are the original fine-tuned models without compression. We also compare with 8-bit RTN and 4-bit GPTQ (Frantar et al., 2022) on evaluations where we run BitDelta on quantized base models.
Models and datasets.
We benchmark fine-tuned models based on the Llama-2 (Touvron et al., 2023) and Mistral (Jiang et al., 2023) model families: Vicuna, Xwin-LM, Solar-70B, Zephyr, OpenChat 3.5, Dolphin 2.2.1, and OpenOrca (Chiang et al., 2023; Team, 2023; Upstage, 2023; Tunstall et al., 2023; Wang et al., 2023; Hartford, 2023; Mukherjee et al., 2023). We evaluate on eight tasks: MT-Bench, 25-shot ARC Challenge, 5-shot BBH, 10-shot HellaSwag, zero-shot TruthfulQA, zero-shot LAMBADA, zero-shot Winogrande, and 5-shot GSM8K (Zheng et al., 2023; Clark et al., 2018; Suzgun et al., 2022; Zellers et al., 2019; Lin et al., 2022; Paperno et al., 2016; Sakaguchi et al., 2019; Cobbe et al., 2021). We use FastChat (Zheng et al., 2023) to evaluate on MT-Bench, and use lm-evaluation-harness (Gao et al., 2023) to evaluate on the other tasks. We denote our methodology before scale distillation is applied as BitDelta-Initial.
We primarily focus on high-margin metrics where fine-tuning is significantly impactful and aggregate the other metrics. See Tables 6 to 9 in the Appendix for full results. BitDelta performs quite well on the aggregated metrics, even outperforming the baseline in many cases. However, it’s important to contextualize these results with regard to the base model itself, which is also performant on these metrics.
In such cases, it’s difficult to attribute performance to our methodology or to the underlying base model. Because of this, we highlight TruthfulQA, GSM8K, and MT-Bench, which base models tend to struggle on, to show that BitDelta accurately preserves fine-tune information.
4.2 Accurate Quantization
SVD comparison.
We compare BitDelta to a low rank approx. of the weight delta on Vicuna-7B v1.5. For the low rank approx., we decompose $\Delta=U\Sigma V$ and approximate $\hat{\Delta}=AB$ where $A=U\sqrt{\hat{\Sigma}}$, $B=\sqrt{\hat{\Sigma}}V$. During distillation, we treat all entries of the low rank matrices as trainable parameters. We find that the low rank approx. fails to fully capture the fine tune information, and underperforms across the board (Table 1). Interestingly, we find that distillation is not as effective for the low rank approx. compared to BitDelta, though we partly attribute this to the excess of trainable parameters leading to training instability.
Main Results.
BitDelta is performant across various model families, across a wide range of model sizes, and across many fine-tuning techniques. We benchmark on Llama-2 and Mistral families, and on models ranging from 7B to 70B parameters. Shown in Table 2, we find that BitDelta is very general and can recover all types of finetune information, including SFT-based methods (Radford et al., 2018) on Mistral-7B v0.1 Instruct, RLHF-based methods (Christiano et al., 2023) on Llama 2 Chat, and context extension methods (RoPE scaling) (Chen et al., 2023c; Press et al., 2022) on Vicuna-7B v1.5 16k.
Figure 3: As the fidelity of $\Delta$ increases, the TruthfulQA scores of Llama 2-7B + $\Delta$ approaches that of Vicuna-7B v1.5.
We note that GSM8K for BitDelta-Initial on Mistral-7B v0.1 Instruct and Zephyr-7B-$\beta$ is abnormally high; we attribute this to how performant the base model Mistral-7B v0.1 is on this task in comparison.
Scale distillation is effective. Introducing scale distillation raises TruthfulQA and GSM8K scores to within 1-2 points of the baseline fine-tune, and generally raises MT-Bench scores to within 0.1-0.2 points.
Case Study.
We present a sample response from Zephyr-7B-$\beta$ in Table 3, highlighting the efficacy of scale distillation. BitDelta-Initial does not have a casual tone, and makes no attempt to adhere to the word limit. With the introduction of scale distillation, BitDelta exhibits greater instruction following capabilities, producing a catchy response that slightly exceeds the word limit.
Quantized base models.
Because 8-bit RTN and GPTQ work with 16-bit activations, we can keep the fine-tune weights $W_{\text{fine}}$ and scaling factors $\alpha$ in high precision, only quantizing the base weights $W_{\text{base}}$. As shown in Table 4, we find that BitDelta is still performant when applied to quantized base models.
We observe that FP16 + $\Delta$ outperforms GPTQ across the board. In the performance engineering context of multi-tenancy serving, as we add more models, we would rather store a single high precision base model with many 1-bit deltas than store many quantized fine-tuned models. This interesting result implies that the above also holds true in the model quality context of multi-tenancy serving.
We try using Llama 2-7B Chat as both the base model and fine-tune model, quantizing the base model using GPTQ, and find that we’re able to outperform baseline GPTQ on many evaluations. We hypothesize this is because we can diffuse 16-bit information into the model through high precision scaling factors, at the cost of including a 1-bit delta.
Ablation over fidelity of $\Delta$.
By successively applying BitDelta, treating the compressed model from the previous iteration as our base model, we can vary the granularity over the delta, associating it with multiple 1-bit masks. One advantage of doing this is the ability to assign arbitrary scale factors to each 1-bit mask. In contrast, when increasing the bit size, scale factors are implicitly fixed with respect to each other. Figure 3 shows how the TruthfulQA of Llama 2-7B plus an increasingly granular delta approaches that of Vicuna-7B v1.5. Full results are in Table 8.
Table 5: BitDelta achieves over 10$\times$ compression. We can further compress the embedding and LM head layers, but leave this to future work due to inconsistencies in tokenizer vocabularies.
Base Model Size $\Delta$ Size Comp. Factor
Llama 2-7B 13.48 GB 1.24 GB 10.87
Llama 2-13B 26.03 GB 2.09 GB 12.45
Llama 2-70B 137.95 GB 8.95 GB 15.41
Mistral-7B v0.1 14.48 GB 1.30 GB 11.14
4.3 Latency Improvement
To illustrate the idea of translating memory saving into improved latency, we implement a simple Triton kernel for GEMM with a binary matrix and scaling factor, as in BitDelta. We assume there are $B$ fine-tuned models served on the same GPU. For simplicity, we consider the setting where the requests are uniformly distributed, i.e., each model receives one request simultaneously. Following the decomposition in Eq. (6), the kernel is used to compute the batched matrix multiplication between $B$ binary matrices ($N\times M$) and $B$ high-precision activations ($L\times N$) where $N,M$ are intermediate dimensions and $L$ is the sequence length. We focus on decoding latency as opposed to the prefill latency, as it consumes most of the time. Tokens are generated one by one in decoding, meaning $L$ is always 1.
Kernel latency.
We benchmark the decoding latency of our kernel, a batched linear operation over multiple deltas with a single base model, as in Eq. (6), corresponding to a single linear layer, and compare against naively computing the forward pass separately for each model. We ablate across the batch size and hidden size dimensions and find that our kernel consistently achieves about 2$\times$ speedup.
Figure 4: Decoding latency of a linear layer with and without BitDelta. Blue: Naive forward pass with $B$ distinct fine-tuned models. Yellow: Batched forward pass with BitDelta, corresponding to one base model and $B$ 1-bit deltas, utilizing a Triton kernel. Left: Ablation over hidden size, assuming $N=M$ and $B=8$. Right: Ablation over batch size, assuming $N=M=8192$.
End-to-end latency.
We also benchmark the end-to-end decoding latency on Llama 2-7B variants with an input length of 128 (we find the decoding latency is less sensitive to the input length), ablated across the batch size. We find that BitDelta introduces overhead when the batch size is low. However, BitDelta scales better and successfully translates the saved GPU memory to improved decoding latency, starting at $B=4$. This is exacerbated at higher batch sizes, where the naive approach succumbs to out-of-memory issues and BitDelta is still performant.
Figure 5: End-to-end decoding latency of Llama 2-7B variants with and without BitDelta. Blue: Naive forward pass with $B$ distinct fine-tuned models. Orange: Projected values for the naive forward pass. Green: Batched forward pass with BitDelta. The naive forward pass succumbs to GPU memory issues at higher batch sizes, whereas BitDelta is still performant.
5 Conclusion and Discussion
We propose BitDelta, a simple yet effective approach for efficiently quantizing the weight delta arising from fine-tuning in LLMs down to 1 bit. BitDelta encodes the sign bits of the weight delta and a per-weight matrix scaling factor, which is calibrated further through distillation. This allows for representing multiple full-parameter fine-tuned models with one base model and multiple 1-bit deltas, enhancing applications in multi-tenancy serving by reducing GPU memory requirements and improving generation latency. BitDelta is fast and accurate, showcasing minimal performance degradation, and opens new avenues for efficient model deployment and resource utilization in machine learning.
Future work.
We mainly focus on the qualitative evaluation of BitDelta, as quality is the main challenge that most quantization methods struggle with. We show BitDelta can maintain the quality of the original fine-tuned models and further illustrate the idea of how to translate the memory saving into improved latency with a fused kernel. Yet, there is significant room for improvement – our kernel is fairly slow compared to the theoretical limit, considering the small memory footprint of binary matrices. It may also be possible to achieve better model quality by preserving salient weight deltas. Finally, the idea of calibrating certain scale factors through distillation may be applied more generally to PTQ methods. We hope our work motivates future refinement of scale factors, making low-bit quantized LLMs more robust.
6 Impact Statement
Environmental Sustainability and Cost Reduction.
The reduction in GPU memory requirements through BitDelta translates to lower energy consumption and a reduction in costs associated with serving multiple fine-tuned models. By significantly lowering hardware requirements, BitDelta contributes to making AI technologies more eco-friendly and economically viable. This approach aligns with the growing need for sustainable and cost-effective computing solutions in the deployment of large-scale AI technologies.
Democratization of Fine-tuned Models.
By dramatically reducing the hardware requirements for serving fine-tuned models, BitDelta enables smaller entities to deploy state-of-the-art models more feasibly. This can accelerate innovation and application development across various industries and academic fields, making fine-tuned models accessible to a wider audience.
Dealignment Mitigation.
BitDelta is a lossy compression method on the fine-tune information in LLMs. As such, crucial alignment information may be lost in the process of compression. We believe this is an important consequence to highlight, as BitDelta democratizes multi-tenant applications which may exacerbate this dealignment concern. We encourage further work on evaluation techniques to detect alignment loss in BitDelta, which can lead to the creation of robust methods for its mitigation.
Acknowledgments
We thank Together AI, MyShell AI, National Science Foundation (NSF), MIT-IBM Watson AI Lab, and MIT Amazon Science Hub for supporting this research.
Appendix A Appendix
Table 6: Full results of the comparison of BitDelta to the SVD based method, corresponding to Table 1.
Model/Method Train Loss ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average normal-$\uparrow$ MT-Bench
Llama 2-7B 52.56 33.76 78.96 38.96 68.39 68.98 13.57 50.74
Vicuna-7B v1.5 53.92 37.14 77.45 50.36 64.41 69.61 19.03 53.13 6.04
BitDelta-Initial 0.41 54.69 36.74 78.47 47.63 66.31 68.75 19.56 53.16 5.67
BitDelta 0.052 54.27 36.57 77.9 49.97 65.2 69.46 20.17 53.36 5.99
SVD-Initial 0.78 54.61 33.96 78.61 44.56 67.77 68.03 15.92 51.92 4.45
SVD 0.34 53.92 34.51 77.95 44.55 66.35 68.27 14.18 51.39 4.98
Table 7: Full results of the application of BitDelta to quantized base models, corresponding to Table 4.
Base Model Method ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average normal-$\uparrow$ MT-Bench
Baseline FP16 53.58 33.84 78.58 45.32 66.58 66.46 22.74 52.44 6.56
LLM.int8() 53.24 33.71 78.62 45.02 66.5 66.06 22.29 52.21 6.28
GPTQ 51.88 33.54 77.17 44.92 65.32 65.43 19.48 51.11 5.90
Llama 2-7B FP16 + $\Delta$ 54.44 33.85 78.31 44.95 66.66 66.14 20.24 52.08 6.47
LLM.int8() + $\Delta$ 53.67 33.48 78.57 44.71 66.7 66.85 19.86 51.98 6.16
GPTQ + $\Delta$ 51.45 33.90 78.06 42.52 66.85 65.82 19.94 51.22 6.02
Llama 2-7B Chat GPTQ + $\Delta$ 52.56 33.65 77.54 44.63 65.81 66.30 22.14 51.80 6.11
Table 8: Full results of the ablation over the fidelity of $\Delta$, corresponding to Figure 3.
# bits in $\Delta$ ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average normal-$\uparrow$
Llama 2-7b 52.56 33.76 78.96 38.96 68.39 68.98 13.57 50.74
1 bit 54.27 36.57 77.90 49.97 65.20 69.46 20.17 53.36
2 bits 54.44 36.78 77.71 49.69 65.26 69.22 20.62 53.39
3 bits 54.27 36.94 77.58 49.90 65.11 70.09 19.48 53.34
4 bits 54.18 36.94 77.54 49.80 64.95 69.53 19.18 53.16
5 bits 53.67 36.78 77.63 50.15 65.22 69.69 18.57 53.10
6 bits 53.67 36.85 77.64 50.20 65.07 69.69 18.80 53.13
7 bits 53.74 37.01 77.56 50.29 65.15 69.38 18.50 53.09
8 bits 53.84 36.94 77.51 50.15 64.95 70.17 18.80 53.19
Vicuna-7b v1.5 53.92 37.14 77.45 50.36 64.41 69.61 19.03 53.13
Table 9: Full results of BitDelta applied to fine-tuned models in the Llama-2 and Mistral families, corresponding to Table 2.
Model Method ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average normal-$\uparrow$ MT-Bench $\uparrow$
Llama 2-7B 52.56 33.76 78.96 38.96 68.39 68.98 13.57 50.74
Llama 2-7B Chat Baseline 53.58 33.84 78.58 45.32 66.58 66.46 22.74 52.44 6.56
BitDelta-Initial 55.46 35.56 76.32 41.10 68.14 68.03 18.27 51.84 6.31
BitDelta 54.44 33.85 78.31 44.95 66.66 66.14 20.24 52.08 6.47
Vicuna-7B v1.5 Baseline 53.92 37.14 77.45 50.36 64.41 69.61 19.03 53.13 6.04
BitDelta-Initial 54.69 36.74 78.47 47.63 66.31 68.75 19.56 53.16 5.67
BitDelta 54.27 36.57 77.9 49.97 65.2 69.46 20.17 53.36 5.99
Vicuna-7B v1.5 16k Baseline 54.86 35.63 77.06 50.38 52.32 67.64 14.18 50.30 6.06
BitDelta-Initial 55.55 33.24 77.99 45.58 56.8 68.98 13.95 50.30 5.69
BitDelta 54.61 34.68 77.14 48.75 53.89 67.88 14.48 50.20 6.24
Xwin LM-7B v0.1 Baseline 57.59 34.05 79.15 48.06 68.02 69.22 10.77 52.41 6.24
BitDelta-Initial 56.40 33.90 80.26 44.56 69.86 69.14 16.68 52.97 5.79
BitDelta 57.94 34.19 79.36 47.62 68.29 69.53 9.02 52.28 6.50
Llama 2-13B 59.47 39.03 82.23 36.90 70.44 72.22 22.74 54.72
Llama 2-13B Chat Baseline 60.32 37.89 82.15 43.95 68.62 70.96 33.13 56.72 6.98
BitDelta-Initial 59.90 38.04 82.13 41.70 69.82 71.35 33.36 56.61 7.06
BitDelta 59.98 38.03 81.92 43.47 68.46 71.43 31.92 56.46 6.95
Vicuna-13B v1.5 Baseline 57.34 39.47 81.14 50.86 68.48 71.67 29.72 56.95 6.48
BitDelta-Initial 54.69 36.74 78.47 47.63 66.31 68.75 31.84 54.92 6.51
BitDelta 57.42 39.20 81.33 50.39 68.81 71.51 30.48 57.02 6.81
Vicuna-13B v1.5 16k Baseline 54.86 35.63 77.06 50.38 52.32 67.64 29.72 52.52 6.90
BitDelta-Initial 59.90 38.04 82.13 41.70 69.82 71.35 26.76 55.67 6.60
BitDelta 54.61 34.68 77.14 48.75 53.89 67.88 28.73 52.24 6.88
WizardLM-13B v1.2 Baseline 60.15 40.82 82.58 47.17 69.26 71.90 42.38 59.18 6.95
BitDelta-Initial 60.41 40.27 83.26 44.89 70.23 71.74 42.08 58.98 6.73
BitDelta 60.92 41.30 82.55 46.67 68.97 71.51 41.62 59.08 6.93
Xwin LM-13B v0.1 Baseline 63.14 40.12 82.92 45.54 70.62 73.09 21.15 56.65 6.78
BitDelta-Initial 63.4 40.33 83.71 43.6 71.26 73.09 26.76 57.45 6.70
BitDelta 62.80 39.81 83.01 48.19 70.74 72.30 21.76 56.94 6.83
Llama 2-70B 67.58 51.67 87.00 44.82 74.81 77.98 52.69 65.22
Llama 2-70B Chat Baseline 65.44 43.93 85.91 52.77 73.90 74.90 47.61 63.49 7.12
BitDelta-Initial 63.4 38.67 81.36 41.63 72.66 73.95 42.38 59.15 6.85
BitDelta 65.87 44.97 85.65 51.37 74.29 74.90 48.82 63.70 7.06
Solar-0-70B Baseline 71.16 55.54 87.78 62.03 75.04 79.32 56.18 69.58 7.07
BitDelta-Initial 69.54 54.52 87.57 59.08 75.37 78.69 56.79 68.79 6.79
BitDelta 70.82 55.06 87.35 62.03 75.86 78.77 56.63 69.50 6.82
Xwin LM-70B v0.1 Baseline 70.65 52.40 87.15 60.06 75.04 78.06 40.33 66.24 7.45
BitDelta-Initial 69.97 52.93 87.36 60.77 75.51 78.14 50.64 67.90 7.70
BitDelta 70.22 52.22 86.97 58.57 75.49 77.58 40.18 65.89 7.34
Mistral-7B v0.1 61.35 41.18 83.46 42.60 70.10 73.80 37.76 58.61
Mistral-7B v0.1 Instruct Baseline 55.03 38.66 75.52 55.93 63.28 69.30 32.75 55.78 6.86
BitDelta-Initial 59.22 40.25 79.91 51.27 67.63 72.14 38.82 58.46 6.54
BitDelta 55.38 37.95 75.62 55.23 66.06 70.48 31.54 56.04 6.43
Zephyr-7B-$\beta$ Baseline 63.82 39.04 84.33 55.12 66.23 72.69 34.34 59.37 7.18
BitDelta-Initial 63.57 41.87 83.85 54.53 67.73 73.56 40.26 60.77 6.70
BitDelta 65.02 41.64 84.05 58.39 66.33 73.95 31.92 60.19 7.00
OpenChat 3.5 Baseline 64.51 45.28 84.39 47.34 65.19 72.61 68.84 64.02 7.74
BitDelta-Initial 64.16 45.23 84.13 43.34 68.62 77.43 57.77 62.95 5.71
BitDelta 64.93 44.57 84.44 46.24 65.88 76.40 57.70 62.88 7.38
Dolphin 2.2.1 Baseline 64.16 44.49 83.30 54.02 69.36 75.22 54.28 63.55 7.36
BitDelta-Initial 64.16 44.43 84.01 48.14 69.98 75.30 50.27 62.33 7.10
BitDelta 64.59 43.08 83.44 54.91 68.39 75.37 52.84 63.23 7.20
OpenOrca-7B Baseline 62.80 44.45 83.58 52.30 66.10 73.24 50.11 61.80 6.70
BitDelta-Initial 63.74 44.46 84.15 49.66 69.05 74.03 49.96 62.15 7.12
BitDelta 63.65 43.46 83.49 51.67 66.12 74.27 49.58 61.75 7.05

### v3

@@
 Large Language Models (LLMs) are typically trained in two phases: pre-training on large internet-scale datasets, and fine-tuning for downstream tasks.
-Given the higher computational demand of pre-training, it’s intuitive to assume that fine-tuning adds less new information to the model, and is thus more compressible.
+Given the higher computational demand of pre-training, it is intuitive to assume that fine-tuning adds less new information to the model, and is thus more compressible.
 We explore this assumption by decomposing the weights of fine-tuned models into their pre-trained components and an additional delta.
-We introduce a simple method, BitDelta, which successfully quantizes this delta down to 1 bit without compromising performance.
+We introduce a simple post-fine-tuning method, BitDelta, which successfully quantizes this delta down to 1 bit without compromising performance.
 This interesting finding not only highlights the potential redundancy of information added during fine-tuning, but also has significant implications for the multi-tenant serving and multi-tenant storage of fine-tuned models.
-By enabling the use of a single high-precision base model accompanied by multiple 1-bit deltas, BitDelta dramatically reduces GPU memory requirements by more than 10$\times$, which can also be translated to enhanced generation latency in multi-tenant settings.
-We validate BitDelta through experiments across Llama-2 and Mistral model families, and on models up to 70B parameters, showcasing minimal performance degradation over all tested settings.
-Code is avaliable at https://github.com/FasterDecoding/BitDelta.
-Machine Learning, ICML
-Machine Learning, ICML
+By enabling the use of a single high-precision base model accompanied by multiple 1-bit deltas, BitDelta dramatically reduces GPU memory requirements by more than 10$\times$, thus reducing per-user generation latency by more than 10×10$\times$10 × in multi-tenant settings.
+We validate BitDelta through experiments across Llama-2, Mistral and MPT model families, and on models up to 70B parameters, showcasing minimal performance degradation in all tested settings.
 1 Introduction
-After large-scale pretraining, foundation models are typically fine-tuned for specific downstream tasks (Devlin et al., 2019; Radford et al., 2018, 2019).
-This pretrain-finetune paradigm has revolutionized machine learning; LLMs have not only proven effective for critical tasks such as instruction following and alignment (Ouyang et al., 2022), but are also performant on a wide array of niche yet highly impactful applications (Xu et al., 2024; Qiu et al., 2023).
+After large-scale pretraining, foundation models are typically fine-tuned for specific downstream tasks [16, 43, 44].
+This pretrain-finetune paradigm has revolutionized machine learning; LLMs have not only proven effective for critical tasks such as instruction following and alignment [39], but are also performant on a wide array of niche yet highly impactful applications [61, 42].
 Through fine-tuning, LLMs are adeptly equipped to align with distinct user preferences or specialized task requirements, showcasing an unprecedented level of adaptability.
@@
 By compressing this delta while maintaining model performance, we aim to sidestep the prohibitive costs associated with storage and GPU memory demands.
-From the delta decomposition point of view, parameter-efficient fine-tuning (PEFT) methods like LoRA (Hu et al., 2021; Houlsby et al., 2019a; Rebuffi et al., 2017; Dettmers et al., 2023; Chen et al., 2023d) effectively enforce a highly structured and compressed form of delta during fine-tuning, a powerful insight for model serving of PEFT-based fine-tunes.
-(Sheng et al., 2023) and (Chen et al., 2023b) explore multi-tenant serving of LoRA-based fine-tunes.
+From the delta decomposition point of view, parameter-efficient fine-tuning (PEFT) methods like LoRA [25, 24, 46, 15, 9] effectively enforce a highly structured and compressed form of delta during fine-tuning, a powerful insight for model serving of PEFT-based fine-tunes.
+Sheng et al.
+[49] and Chen et al.
+[7] explore multi-tenant serving of LoRA-based fine-tunes.
 Figure 1: Overview of BitDelta.
@@
 BitDelta shows minimal degradation in model performance and reduces memory consumption in multi-tenancy serving by representing multiple fine-tuned models with a single high-precision base model and multiple 1-bit deltas.
-Nevertheless, recent work has shown that PEFT methods may not yet match the model quality of full parameter fine-tuning, especially on high resource tasks (Chen et al., 2022), and are fairly sensitive to hyperparameter choice and prompting methods (Niederfahrenhorst et al., 2023).
-As a result, we notice that among the 2307 LLMs (as of time of writing) on the Open LLM Leaderboard (Beeching et al., 2023) with a valid README file, only $<20\%$ indicate that they exclusively use LoRA.
-Most models are full parameter fine-tunes, model merges (Yu et al., 2023; Jin et al., 2023; Wortsman et al., 2022) of full parameter fine-tunes, or model merges of LoRA based fine-tunes (which are effectively high-rank).
-It is also attractive to approximate general deltas with low-rank matrices post-training.
+Nevertheless, recent work has shown that PEFT methods may not yet match the model quality of full parameter fine-tuning, especially on high resource tasks [6], and are fairly sensitive to hyperparameter choice and prompting methods [38].
+Biderman et al.
+[2] show that LoRA’s reduced expressivity, although providing desirable regularization, leads to significantly worse performance compared to full fine-tuning in math and programming tasks.
+As a result, we notice that among the 2307 LLMs (as of time of writing) on the Open LLM Leaderboard [1] with a valid README file, only $<20\%$ indicate that they exclusively use LoRA.
+Most models are full parameter fine-tunes, model merges [64, 28, 59] of full parameter fine-tunes, or model merges of LoRA based fine-tunes (which are effectively high-rank).
+Figure 2: Cumulative Explained Variance (CEV) plot of a $4096\times 4096$ weight delta between Llama 2-7B and Vicuna-7B v1.5.
+Deltas from full parameter fine-tuning are fairly high rank, making low-rank approximations difficult.
+It is also attractive to approximate general deltas with low-rank matrices post-training (in particular, post-fine-tuning).
 However, experimental results show that this is challenging (Table 1), as deltas from full parameter fine-tunes tend to be fairly high-rank (Figure 2).
-Figure 2: CEV plot of a $4096\times 4096$ weight delta between Llama 2-7B and Vicuna-7B v1.5. Deltas from full parameter fine-tuning are fairly high rank, making low-rank approximations difficult.
-Instead, we draw from the insight that motivates PEFT methods in general: Given the higher computational demand of pre-training, it’s intuitive to assume that fine-tuning adds less new information to the model, and is thus much more compressible.
-Indeed, we find that we can efficiently quantize the delta to merely 1 bit with almost no performance drop.
+We instead draw from the insight that motivates PEFT methods in general: Given the higher computational demand of pre-training, it is intuitive to assume that fine-tuning adds less new information to the model, and is thus much more compressible.
+In fact, we find that we can efficiently quantize the delta to merely 1 bit with almost no performance drop.
 We propose BitDelta, an efficient post-training quantization (PTQ) solution that acts on the weight delta between a fine-tuned model and its underlying base model.
-BitDelta has two stages: 1) We quantize the delta of each weight matrix into a scaling factor multiplied by a binary matrix.
-Specifically, we take the sign of the delta to form the binary matrix and initialize the scaling factor as the average of the absolute values of the delta, which minimizes $L_{2}$ norm.
+Table 1: Comparison between BitDelta and a SVD based method, with Llama 2-7B and Llama 2-7B Chat as the base and fine-tuned models.
+BitDelta is performant across the board, whereas the SVD-based method fails to sufficiently capture the fine-tuned information.
+Model/Method TruthfulQA GSM8K MT-Bench Adjusted Average1 $\uparrow$
+Llama 2-7B 38.96 13.57 60.53
+Llama 2-7B Chat 45.32 22.74 6.56 59.81
+BitDelta-Initial 41.10 18.27 6.31 60.70
+BitDelta 44.95 20.24 6.47 59.88
+SVD-Initial ($r=16$) 42.57 7.13 4.73 60.58
+SVD ($r=16$) 42.42 5.05 4.99 60.71
+SVD-Initial ($r=128$) 43.90 17.82 5.68 60.21
+SVD ($r=128$) 43.32 11.83 5.85 60.58
+BitDelta consists of two stages: 1) We quantize the delta between a fine-tuned model’s weight matrix and base model’s weight matrix into a scaling factor multiplied by a binary matrix.
+Specifically, we take the sign of the weight delta to form the binary matrix and initialize the scaling factor as the average of the absolute values of the delta, minimizing $L_{2}$ quantization error.
 2) We further calibrate the scaling factors through model distillation over a small calibration dataset while keeping the binary matrices frozen.
-Despite the small number of trainable parameters and training steps, we find that this distillation process is effective in further recovering model quality.
-BitDelta is extremely efficient compared to other quantization methods; we are able to compress 70B models in roughly 10 minutes, whereas methods like GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2023) may take multiple GPU hours.
-Our experiments over 17 popular fine-tuned models affirm that BitDelta can be applied across various model types and sizes with minimal impact on performance.
-BitDelta opens up opportunities to efficiently serve multiple fine-tuned models with shared servers: By only storing a single full-precision base model, and (dynamically) loading and performing batched inference over multiple 1-bit deltas, we can efficiently represent multiple fine-tuned models.
-Compared to naively using full precision fine-tuned models, deltas compressed by BitDelta are more than 10$\times$ smaller, consuming less GPU memory, and can therefore be loaded faster.
+Despite the small number of trainable parameters and calibration steps, we find that this distillation process is effective in further recovering model quality.
+Our experiments over 17 popular fine-tuned models affirm that BitDelta can be applied across various model types and model sizes with minimal impact on performance.
+BitDelta creates opportunities to efficiently serve multiple fine-tuned models with shared servers: By only storing a single full-precision base model, and (dynamically) loading and performing batched inference over multiple 1-bit deltas, we can efficiently represent multiple fine-tuned models.
+Compared to naively using full precision fine-tuned models, deltas compressed by BitDelta are more than 10$\times$ smaller, and can therefore be loaded faster.
 This addresses the storage challenge.
-Moreover, since LLM inference is memory-bound (Leviathan et al., 2022; Chen et al., 2023a; Cai et al., 2024), the latency of each decoding step is proportional to the GPU memory consumption of the model weights.
-With an efficient CUDA kernel implementation, we can translate this memory reduction into a latency speedup, similar to other quantization methods (Frantar et al., 2022; Lin et al., 2023).
-As a preliminary attempt, we implement a 1-bit matrix multiplication kernel in Triton (Tillet et al., 2019) and show that it can improve the multi-tenant serving latency by about 2$\times$, addressing the serving challenge.
-Finally, we study a few extensions of BitDelta, where we quantize the base model, and where we iteratively apply BitDelta.
+Moreover, since LLM inference is memory-bound [32, 5, 3], the latency of each decoding step is proportional to the GPU memory consumption of the model weights.
+With an efficient CUDA kernel implementation, we can translate this memory reduction into a latency reduction, similar to other quantization methods [19, 33].
+Using the $W_{INT1}A_{FP16}$ kernel from BitBLAS [58], we improve the multi-tenant serving latency of full-parameter fine-tuned models by more than 10×10$\times$10 ×.
+Finally, we study a few extensions of BitDelta, where we quantize the base model and where we iteratively apply BitDelta.
 Experimental results show that our method is quite general and can be applied to various use cases.
-Table 1: Comparison between BitDelta and a SVD based method ($r=16$), with Llama 2-7B and Vicuna-7B v1.5 as the base and fine-tuned models.
-BitDelta is performant across the board, whereas the SVD-based method fails to sufficiently capture the fine-tuned information.
-Model/Method Train Loss TruthfulQA GSM8K MT-Bench Adjusted Average2.1 $\uparrow$
-Llama 2-7B 38.96 13.57 60.53
-Vicuna-7B v1.5 50.36 19.03 6.04 60.51
-BitDelta-Initial 0.41 47.63 19.56 5.67 60.99
-BitDelta 0.052 49.97 20.17 5.99 60.68
-SVD-Initial 0.78 44.56 15.92 4.45 60.60
-SVD 0.34 44.55 14.18 4.98 60.20
+22footnotetext: Adjusted Average is over ARC, BBH, HellaSwag, WinoGrande, and excludes TruthfulQA, GSM8K, MT-Bench.
 2 Related Work
@@
 Xiao et al.
-(2023) implement a technique that rescales between activations and parameters, effectively mitigating outlier activations to facilitate smoother quantization.
+[60] implement a technique that rescales between activations and parameters, effectively mitigating outlier activations to facilitate smoother quantization.
 Dettmers et al.
-(2022) develop an approach that decomposes matrix multiplications into 8-bit computations, with an additional 16-bit process for handling outliers.
+[14] develop an approach that decomposes matrix multiplications into 8-bit computations, with an additional 16-bit process for handling outliers.
 Exploring further, Frantar et al.
-(2022) introduce a method that iteratively rounds weight columns to 3-4 bits of precision.
+[19] introduce a method that iteratively rounds weight columns to 3-4 bits of precision.
 Similarly, Lin et al.
-(2023) propose an activation-aware quantization scheme that selectively preserves crucial weights while compressing the majority to 3-4 bits.
-In a different vein, Kim et al.
-(2023) devise a sparse, low-precision pattern focusing on a small yet significant set of weights.
-Lastly, Chee et al.
-(2023) utilize incoherence processing to quantize model weights to as low as 2 bits with minimal impact on performance.
+[33] propose an activation-aware quantization scheme that selectively preserves crucial weights while compressing the majority to 3-4 bits.
+Kim et al.
+[29] devise a sparse, low-precision pattern focusing on a small yet significant set of weights.
+Chee et al.
+[4] utilize incoherence processing to quantize model weights to as low as 2 bits with minimal impact on performance.
 Pruning.
 Pruning also aims to reduce the memory consumption of neural networks.
-It accomplishes this by pushing certain parameter values to zero, inducing sparsity in the model (LeCun et al., 1989; Han et al., 2015, 2016; Zhu & Gupta, 2017).
-However, these methods may fail to take advantage of modern hardware like GPUs unless using certain structured sparsity patterns like 2:4 (50%) sparsity (Mishra et al., 2021).
-Frantar & Alistarh (2023) demonstrate a pruning method on LLMs that successfully utilizes the 2:4 sparsity pattern and achieves a 50% sparsity ratio.
+It accomplishes this by pushing certain parameter values to zero, inducing sparsity in the model [31, 21, 22, 67].
+However, these methods may fail to take advantage of modern hardware like GPUs unless using certain structured sparsity patterns like 2:4 (50%) sparsity [36].
+Frantar and Alistarh [18] demonstrate a pruning method on LLMs that successfully utilizes the 2:4 sparsity pattern and achieves a 50% sparsity ratio.
 It is challenging to obtain higher sparsity while being hardware-friendly.
-22footnotetext: Adjusted Average is over ARC, BBH, HellaSwag, WinoGrande, and excludes TruthfulQA, GSM8K, MT-Bench.
-2.2 Delta Compression
-Parameter-efficient fine-tuning.
-Parameter-efficient fine-tuning (PEFT) techniques reduce the number of trainable parameters during fine-tuning, reducing memory and compute demand while achieving promising accuracy.
-Houlsby et al.
-(2019b) introduce adapter layers, freezing the base model during fine-tuning.
-Hu et al.
-(2021) propose Low-Rank Adaptation (LoRA), which enforces that weight updates during fine-tuning are low rank.
-By enforcing structure in the fine-tuning information, PEFT techniques can viewed as compression-aware training.
-Similar to our method, such compression induced by PEFT techniques can also enhance multi-tenant serving.
-Chen et al.
-(2023b) and Sheng et al.
-(2023) explore this idea by developing scalable multi-tenant systems for LoRA-based fine-tunes, based on customized CUDA kernels that exploit the low-rank structure of the fine-tuned weight deltas.
 Early work on post-training delta compression.
-Most related to our work, a few studies explore the idea of post-training delta compression by adopting existing compression techniques like GPTQ, unstructured pruning (Han et al., 2016), or even classic lossless compression algorithms.
+Most related to our work, a few studies explore the idea of post-training delta compression by adopting existing compression techniques like GPTQ, unstructured pruning [22], or even classic lossless compression algorithms.
 Isik et al.
-(2023) focus on reducing the delta size to save storage.
+[26] focus on reducing the delta size to save storage.
 Yu et al.
-(2023) utilize pruning to improve model merging applications.
+[64] utilize pruning to improve model merging applications.
 Yadav et al.
-(2023) reduces the size of PEFT modules to save storage.
+[62] reduces the size of PEFT modules to save storage.
 Ryu et al.
-(2023) combines quantization with low-rank approximation to reduce the delta size.
-The concurrent and independent work by Yao & Klimovic (2023) also explores using delta compression to improve multi-tenant serving, but focuses more on reducing the model loading time from disk to GPU.
-Compared to existing work, we offer a much simpler and faster method, BitDelta, achieving a more than 10$\times$ compression ratio while also being friendly to modern accelerators.
+[47] combines quantization with a low-rank approximation to reduce the delta size.
+The concurrent and independent work by Yao and Klimovic [63] also explores using delta compression to improve multi-tenant serving, but focuses more on reducing the model loading time from disk to GPU.
+Compared to existing work, we offer a much simpler and faster method, BitDelta, achieving a compression ratio of more than 10$\times$ while also being friendly to modern accelerators.
 3 BitDelta
 3.1 Method
-BitDelta consists of two stages: 1) We quantize each weight matrix into a scalar multiplied by a binary matrix***In our experiments, we only quantize the linear layers in the Transformer blocks as they contribute most of the parameters and computation..
-2) We further calibrate the scalar factors with distillation.
+BitDelta consists of two stages: 1) We quantize each weight matrix into a scalar multiplied by a binary matrix***In our experiments, we only quantize the linear layers in the Transformer blocks as they contribute the majority of the parameters and computation..
+2) We further calibrate the scalar factors using model distillation.
 We describe each stage in this section:
 1-bit quantization.
-Let $W_{\text{base}},W_{\text{fine}}\in\mathbb{R}^{n\times m}$ be weight matrices from the base model and fine-tuned model, respectively.
+Let $W_{\text{base}},W_{\text{fine}}\in\mathbb{R}^{n\times m}$ be weight matrices from the base model and fine-tuned model respectively.
 We define the weight delta as $\Delta=W_{\text{fine}}-W_{\text{base}}$, representing the modification in weights post-fine-tuning.
-For efficient representation, we aim to obtain a binarized estimator of this weight delta, denoted as $\hat{\Delta}$, by encoding its sign bits:
+For efficient representation of this weight delta, we aim to obtain a binarized estimator by encoding its sign bits, denoted as $\hat{\Delta}$:
 $\hat{\Delta}=\alpha\odot\text{Sign}(\Delta),$ (1)
@@
 $\text{Sign}(W_{ij})=\begin{cases}+1,&\text{if }W_{ij}>0,\\ -1,&\text{if }W_{ij}\leq 0,\end{cases}$ (2)
-and $\alpha$ is a high-precision scaling factor for the entire matrix. To minimize the quantization error in $L_{2}$ norm:
-$\left\|\Delta-\hat{\Delta}\right\|_{2}^{2}=\sum_{ij}\left(\left|W_{ij}\right|-\alpha\right)^{2},$ (3)
-we take
-$\alpha=\frac{1}{nm}\sum_{ij}|W_{ij}|.$ (4)
+and $\alpha$ is a high-precision scaling factor for the entire matrix. To minimize the quantization error of $\Delta$ in $L_{2}$ norm:
+$\left\lVert\Delta-\hat{\Delta}\right\rVert_{2}^{2}=\sum_{ij}(|W_{ij}|-\alpha)^{2},$ (3)
+we initialize $\alpha$ as follows:
+$\alpha=\frac{1}{nm}\sum_{ij}|\Delta_{ij}|.$ (4)
 Surprisingly, we find that the above quantization approach already does quite well and retains most of the fine-tuned models’ performance.
@@
 BitDelta works for many types of fine-tuned information, including SFT-based methods, RLHF-based methods, and context extension methods (RoPE scaling).
-Scale distillation is effective; it raises TruthfulQA/GSM8K scores to within 1-2 points of the baseline fine-tune, and MT-Bench scores to within 0.1-0.2 points.
-Model Method TruthfulQA GSM8K MT-Bench Adjusted Average2.1 $\uparrow$
+Scale distillation is effective, raising TruthfulQA/GSM8K scores to within 1-2 points of the baseline fine-tune, and MT-Bench scores to within 0.1-0.2 points.
+Model Method TruthfulQA GSM8K MT-Bench Adjusted Average1 $\uparrow$
 Llama 2-7B 38.96 13.57 60.53
@@
 BitDelta 46.67 41.62 6.93 61.86
+Scale distillation.
+The scaling factor $\alpha$ intuitively plays a more significant role in the low-bit regime.
+Additionally, per-matrix $L_{2}$ weight error is not a perfect measure of degradation in overall model quality.
+We further optimize these scales by performing model distillation to align the output logits of the quantized model to that of the original fine-tuned model.
+More concretely, we freeze the model weights and optimize for the following objective:
+$\boldsymbol{\alpha}^{*}=\arg\min_{\boldsymbol{\alpha}}\mathbb{E}_{x\sim\mathbf{X}}\left[\left\|\mathbf{Z}_{\text{fine}}(x)-\mathbf{Z}_{\text{bin}}(x;\boldsymbol{\alpha})\right\|^{2}\right]$ (5)
+where $\mathbf{X}$ is a calibration dataset, and $\mathbf{Z}(\cdot)$ are the logits of the respective models.
+Scale distillation is fairly robust to choice $\mathbf{X}$, as 1) the process is extremely parameter efficient, and 2) the crucial aspect of the process is to logit match with the fine-tuned model, regardless of the actual text content.
+For our experiments, we distill on the C4 dataset [45], consisting of generic internet data, using 800 samples of length 128.
+We use the same subset of C4 over all models to control for seed-based variations.
+We use the Adam optimizer [30] with $lr=10^{-4}$, $\beta=(0.9,0.999)$, $\epsilon=10^{-8}$.
+1x80 GB A100 GPU is used to distill 7B and 13B models, and 6x80GB A100 GPUs are used to distill 70B models (2x for finetune, 4x for binarized).
+Scale distillation is fast; we can compress 70B models in roughly 10 minutes.
+3.2 Methodology Cost
+Compared to full parameter and parameter efficient fine-tuning methods, BitDelta is extremely cheap.
+While fine-tuning methods require training thousands to millions of parameters, BitDelta only necessitates training a single parameter per weight matrix.
+Moreover, BitDelta operates efficiently with input sequences of length 128, unlike fine-tuning methods that demand longer sequences to saturate the context window (4k, 8k, etc.).
+Crucially, BitDelta requires only 200 training steps (assuming a batch size of 4), which is significantly less compared to the 10000-1000000 steps at higher batch sizes needed by fine-tuning methods.
+Thus, in terms of methodology cost, we liken BitDelta more to post-training quantization (PTQ) schemes like GPTQ [19] and AWQ [33], rather than full parameter or parameter efficient fine-tuning, while being faster than most PTQ schemes.
+Table 3: Continuation of Table 2.
+Model Method TruthfulQA GSM8K MT-Bench Adjusted Average1 $\uparrow$
 Llama 2-70B 44.82 52.69 71.81
@@
 BitDelta 54.91 52.84 7.20 66.97
-Table 3: Comparison of model responses from Zephyr-7B-$\beta$ for Question 9 in MT-Bench, a concise advertisement task.
-BitDelta-Initial is unable to follow the instructions, producing an advertisement that is overly formal and makes no attempt to adhere to the word limit.
-With the addition of scale distillation, BitDelta successfully produces a concise, catchy advertisement slightly over the word limit.
-*Prompt slightly modified for clarity.
-Prompt
-Please create a slogan about bioenergy that satisfies the following: 1. Make the tone sound casual 2. Embed an advertisement for a company called ”FlexPower” 3. Fewer than 10 words.
-BitDelta
-”FlexPower’s groundbreaking bioenergy solution: Produce fuel without compromising food production!” (10 words)
-This headline is catchy, casual, and highlights the unique selling point of the…
-GPT-4 Score: 8
-BitDelta-Initial
-”FlexPower’s groundbreaking technology unlocks the key to harnessing the power of renewable bioenergy while minimizing ethical concerns”
-GPT-4 Score: 4
-Scale distillation.
-Intuitively, the scaling factor $\alpha$ plays a more significant role in the low-bit regime.
-Additionally, per-matrix $L_{2}$ weight error is not a perfect measure of degradation in overall model quality.
-Thus, we further optimize these scales by performing model distillation to align the output logits of the quantized model to that of the original fine-tuned model.
-More concretely, we freeze the model weights and optimize for the following objective:
-$\bm{\alpha}^{*}=\arg\min_{\bm{\alpha}}\mathbb{E}_{x\sim\mathbf{X}}\left[\left\|\mathbf{Z}_{\text{fine}}(x)-\mathbf{Z}_{\text{bin}}(x;\bm{\alpha})\right\|^{2}\right]$ (5)
-where $\mathbf{X}$ is a calibration dataset, and $\mathbf{Z}(\cdot)$ are the logits of the respective models.
-We find that scale distillation is fairly insensitive to choice $\mathbf{X}$, as 1) the process is extremely parameter efficient, and 2) the crucial aspect of the process is to logit match with the fine-tuned model, regardless of the actual text content.
-For our experiments, we distill on the C4 dataset (Raffel et al., 2023), which is widely used for pre-training, using 800 samples of length 128, and use the Adam optimizer (Kingma & Ba, 2017) with $lr=10^{-4}$, $\beta=(0.9,0.999)$, $\epsilon=10^{-8}$.
-1x80 GB A100 GPU is used to distill 7B and 13B models, and 6x80GB A100 GPUs are used to distill 70B models (2x for finetune, 4x for binarized).
-Sale distillation is fast; we can compress 70B models in roughly 10 minutes, whereas methods like GPTQ (Frantar et al., 2022) and AWQ (Lin et al., 2023) take multiple hours.
+MPT-7B 33.37 6.22 57.95
+MPT 7B-Chat Baseline 40.22 7.96 5.00 56.5
+BitDelta-Initial 38.96 10.01 4.39 57.11
+BitDelta 39.87 8.11 4.94 56.52
 3.3 Implication
-The ability to compress the delta to merely 1-bit opens up multiple opportunities for improving efficiency, enabling more effective model storage (Isik et al., 2023) – where a single base model can be maintained alongside multiple compressed deltas – and facilitating model hot-swapping  (Chen et al., 2023b; Sheng et al., 2023).
+The ability to compress the delta to merely 1-bit opens up multiple opportunities for improving efficiency, enabling more effective model storage [26] – where a single base model can be maintained alongside multiple compressed deltas – and facilitating model hot-swapping  [7, 49].
 With hot-swapping, the base model remains in GPU memory, and compressed deltas are dynamically loaded in accordance to incoming requests.
 In both cases, the compression ratio can be directly translated into reductions in storage needs and loading times.
-Moreover, BitDelta enables the possibility of a multi-tenant serving system like Punica (Chen et al., 2023b) or S-LoRA (Sheng et al., 2023) but for general fine-tuned models instead of just LoRA models.
+Moreover, BitDelta enables the possibility of a multi-tenant serving system like Punica [7] or S-LoRA [49] but for general fine-tuned models instead of just LoRA models.
 Concretely, we consider the scenario where multiple models fine-tuned from the same base model are served with the same server.
@@
 Since LLM inference follows the memory-bound computation pattern where the generation latency is proportional to the GPU memory used by the model weights, the lower memory consumption also suggests the opportunity to improve the serving latency.
-For example, Punica and S-LoRA exploit LoRA’s structure and memory saving by developing a CUDA kernel that can efficiently calculate the batched delta-activation product for low rank deltas.
+For example, Punica and S-LoRA exploit LoRA’s structure and memory saving by computing the activation product between the shared base weight, and low-rank fine-tuned delta weights separately.
 Similarly, we decompose the forward pass of each linear layer as follows:
 $X^{\prime}_{i}=W_{\text{fine},i}X_{i}\approx W_{\text{base}}X_{i}+\underbrace{\hat{\Delta}_{i}X_{i}}_{\text{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{\text{Kernel}}}}$ (6)
-where $X_{i}$ and $X_{i}^{\prime}$ represent input and output features to the $i$-th fine-tuned model, and the base model weight and the delta are computed separately.
+where $X_{i}$ and $X_{i}^{\prime}$ represent input and output features to the $i$-th fine-tuned model, and the base model weight and the 1-bit delta are computed separately.
 For a batch of requests, $W_{\text{base}}X_{i}$ can be computed with the classic batched GEMM kernel.
-We implement a fused binary GEMM kernel in Triton (Tillet et al., 2019) that allows us to calculate $\hat{\Delta}_{i}X$ in a batched setting while keeping the 1-bit deltas quantized until they are transferred to the GPU cache.
+We utilize the BitBLAS  [58] $W_{INT1}A_{FP16}$ kernel that allows us to calculate $\hat{\Delta}_{i}X$ in a batched setting while keeping the 1-bit deltas quantized until they are transferred to the GPU cache.
 This kernel fuses the dequantization operation with the GEMM calculation, reducing the data moving overhead by a large factor.
-Table 4: We apply BitDelta to Llama 2-7B Chat, and find it holds up when the underlying base model is quantized at various levels.
-FP16 + $\Delta$ outperforms baseline GPTQ across the board, implying that in terms of model quality, we would rather store a single high-precision base model with many 1-bit deltas than store many quantized fine-tuned models.
-GPTQ + $\Delta$ with Llama 2-7B Chat as the base model also outperforms baseline GPTQ on many evaluations, because the delta diffuses 16-bit information through high precision scaling factors.
-Base Model Method TruthfulQA GSM8K MT-Bench Adjusted Average2.1 $\uparrow$
-Baseline FP16 45.32 22.74 6.56 59.81
-INT8 RTN 45.02 22.29 6.28 59.63
-GPTQ 44.92 19.48 5.90 58.67
-Llama 2-7B FP16 + $\Delta$ 44.95 20.24 6.47 59.88
-INT8 RTN + $\Delta$ 44.71 19.86 6.16 59.85
-GPTQ + $\Delta$ 42.52 19.94 6.02 59.22
-Llama 2-7B Chat GPTQ + $\Delta$ 44.63 22.14 6.11 59.17
 4 Experiments
 4.1 Setup
+Table 4: Comparison of model responses from Zephyr-7B-$\beta$ for Question 9 in MT-Bench, a concise advertisement task.
+BitDelta-Initial is unable to follow the instructions, producing an advertisement that is overly formal and makes no attempt to adhere to the word limit.
+With the addition of scale distillation, BitDelta successfully produces a concise, catchy advertisement slightly over the word limit.
+*Prompt slightly modified for clarity.
+Prompt Please create a slogan about bioenergy that satisfies the following: 1. Make the tone sound casual 2. Embed an advertisement for a company called "FlexPower" 3. Fewer than 10 words.
+BitDelta "FlexPower’s groundbreaking bioenergy solution: Produce fuel without compromising food production!" (10 words)
+This headline is catchy, casual, and highlights the unique selling point of the…
+GPT-4 Score: 8
+BitDelta-Initial "FlexPower’s groundbreaking technology unlocks the key to harnessing the power of renewable bioenergy while minimizing ethical concerns"
+GPT-4 Score: 4
 Baselines.
 Our primary baselines are the original fine-tuned models without compression.
-We also compare with 8-bit RTN and 4-bit GPTQ (Frantar et al., 2022) on evaluations where we run BitDelta on quantized base models.
+We also compare with 8-bit RTN, 4-bit GPTQ [19], and 2-bit QuIP# [54] on evaluations where we run BitDelta on quantized base models.
 Models and datasets.
-We benchmark fine-tuned models based on the Llama-2 (Touvron et al., 2023) and Mistral (Jiang et al., 2023) model families: Vicuna, Xwin-LM, Solar-70B, Zephyr, OpenChat 3.5, Dolphin 2.2.1, and OpenOrca (Chiang et al., 2023; Team, 2023; Upstage, 2023; Tunstall et al., 2023; Wang et al., 2023; Hartford, 2023; Mukherjee et al., 2023).
-We evaluate on eight tasks: MT-Bench, 25-shot ARC Challenge, 5-shot BBH, 10-shot HellaSwag, zero-shot TruthfulQA, zero-shot LAMBADA, zero-shot Winogrande, and 5-shot GSM8K (Zheng et al., 2023; Clark et al., 2018; Suzgun et al., 2022; Zellers et al., 2019; Lin et al., 2022; Paperno et al., 2016; Sakaguchi et al., 2019; Cobbe et al., 2021).
-We use FastChat (Zheng et al., 2023) to evaluate on MT-Bench, and use lm-evaluation-harness (Gao et al., 2023) to evaluate on the other tasks.
+We benchmark fine-tuned models based on the Llama-2 [53], Mistral [27], and MPT [51] model families: Vicuna, Xwin-LM, Solar-70B, Zephyr, OpenChat 3.5, Dolphin 2.2.1, and OpenOrca [10, 52, 56, 55, 57, 23, 37].
+We evaluate on eight tasks: MT-Bench, 25-shot ARC Challenge, 5-shot BBH, 10-shot HellaSwag, zero-shot TruthfulQA, zero-shot LAMBADA, zero-shot Winogrande, and 5-shot GSM8K [66, 12, 50, 65, 34, 40, 48, 13].
+We use FastChat [66] to evaluate on MT-Bench, and use lm-evaluation-harness [20] to evaluate on the other tasks.
 We denote our methodology before scale distillation is applied as BitDelta-Initial.
 We primarily focus on high-margin metrics where fine-tuning is significantly impactful and aggregate the other metrics.
-See Tables 6 to 9 in the Appendix for full results.
+See Tables 7 to 10 in the Appendix for full results.
 BitDelta performs quite well on the aggregated metrics, even outperforming the baseline in many cases.
 However, it’s important to contextualize these results with regard to the base model itself, which is also performant on these metrics.
-In such cases, it’s difficult to attribute performance to our methodology or to the underlying base model.
+It’s difficult to attribute performance to our methodology or to the underlying base model in such cases.
 Because of this, we highlight TruthfulQA, GSM8K, and MT-Bench, which base models tend to struggle on, to show that BitDelta accurately preserves fine-tune information.
 4.2 Accurate Quantization
+Table 5: BitDelta achieves over 10$\times$ compression. We can further compress the embedding and LM head layers, but leave this to future work due to inconsistencies in tokenizer vocabularies.
+Base Model Size $\Delta$ Size Comp. Factor
+Llama 2-7B 13.48 GB 1.24 GB 10.87
+Llama 2-13B 26.03 GB 2.09 GB 12.45
+Llama 2-70B 137.95 GB 8.95 GB 15.41
+Mistral-7B v0.1 14.48 GB 1.30 GB 11.14
 SVD comparison.
@@
 During distillation, we treat all entries of the low rank matrices as trainable parameters.
+We compare against two settings: $r=16$ (most commonly used) and $r=128$ (memory equivalence with BitDelta).
 We find that the low rank approx.
 fails to fully capture the fine tune information, and underperforms across the board (Table 1).
-Interestingly, we find that distillation is not as effective for the low rank approx.
-compared to BitDelta, though we partly attribute this to the excess of trainable parameters leading to training instability.
+In particular, the low rank approx.
+heavily underperforms on MT-Bench [10], a difficult multi-turn instruction following dataset fairly indicative of real world performance.
+Interestingly, distillation is not as effective for the low rank approx.
+compared to BitDelta.
 Main Results.
 BitDelta is performant across various model families, across a wide range of model sizes, and across many fine-tuning techniques.
-We benchmark on Llama-2 and Mistral families, and on models ranging from 7B to 70B parameters.
-Shown in Table 2, we find that BitDelta is very general and can recover all types of finetune information, including SFT-based methods (Radford et al., 2018) on Mistral-7B v0.1 Instruct, RLHF-based methods (Christiano et al., 2023) on Llama 2 Chat, and context extension methods (RoPE scaling) (Chen et al., 2023c; Press et al., 2022) on Vicuna-7B v1.5 16k.
-Figure 3: As the fidelity of $\Delta$ increases, the TruthfulQA scores of Llama 2-7B + $\Delta$ approaches that of Vicuna-7B v1.5.
+We benchmark on Llama-2, Mistral, and MPT, families, and on models ranging from 7B to 70B parameters.
+Shown in Table 2, we find that BitDelta is very general and can recover all types of finetune information, including SFT-based methods [43] on Mistral-7B v0.1 Instruct, RLHF-based methods [11] on Llama 2 Chat, and context extension methods (RoPE scaling) [8, 41] on Vicuna-7B v1.5 16k.
 We note that GSM8K for BitDelta-Initial on Mistral-7B v0.1 Instruct and Zephyr-7B-$\beta$ is abnormally high; we attribute this to how performant the base model Mistral-7B v0.1 is on this task in comparison.
-Scale distillation is effective.
-Introducing scale distillation raises TruthfulQA and GSM8K scores to within 1-2 points of the baseline fine-tune, and generally raises MT-Bench scores to within 0.1-0.2 points.
+Scale distillation is effective, raising TruthfulQA and GSM8K scores to within 1-2 points of the baseline fine-tune, and generally raising MT-Bench scores to within 0.1-0.2 points.
+Table 6: We apply BitDelta to Llama 2-7B Chat (with corresponding base model Llama 2-7B), and find it holds up when the underlying base model is quantized at various levels.
+Base Model Method TruthfulQA GSM8K MT-Bench Adjusted Average1 $\uparrow$
+Baseline FP16 45.32 22.74 6.56 59.81
+INT8 RTN 45.02 22.29 6.28 59.63
+GPTQ 44.92 19.48 5.90 58.67
+QuIP# 43.69 10.77 5.37 55.82
+Llama 2-7B FP16 + $\Delta$ 44.95 20.24 6.47 59.88
+INT8 RTN + $\Delta$ 44.71 19.86 6.16 59.85
+GPTQ + $\Delta$ 42.52 19.94 6.02 59.22
+QuIP# + $\Delta$ 42.00 9.72 4.96 57.44
 Case Study.
@@
 Quantized base models.
-Because 8-bit RTN and GPTQ work with 16-bit activations, we can keep the fine-tune weights $W_{\text{fine}}$ and scaling factors $\alpha$ in high precision, only quantizing the base weights $W_{\text{base}}$.
+Because 8-bit RTN, GPTQ, and QuIP# work with 16-bit activations, we can keep the fine-tune weights $W_{\text{fine}}$ and scaling factors $\alpha$ in high precision in the compression process, only quantizing the base weights $W_{\text{base}}$.
 As shown in Table 6, we find that BitDelta is still performant when applied to quantized base models.
-We observe that FP16 + $\Delta$ outperforms GPTQ across the board.
-In the performance engineering context of multi-tenancy serving, as we add more models, we would rather store a single high precision base model with many 1-bit deltas than store many quantized fine-tuned models.
-This interesting result implies that the above also holds true in the model quality context of multi-tenancy serving.
-We try using Llama 2-7B Chat as both the base model and fine-tune model, quantizing the base model using GPTQ, and find that we’re able to outperform baseline GPTQ on many evaluations.
-We hypothesize this is because we can diffuse 16-bit information into the model through high precision scaling factors, at the cost of including a 1-bit delta.
+Figure 3: As the fidelity of $\Delta$ increases, the TruthfulQA scores of Llama 2-7B + $\Delta$ approaches that of Vicuna-7B v1.5.
 Ablation over fidelity of $\Delta$.
@@
 Full results are in Table 9.
-Table 5: BitDelta achieves over 10$\times$ compression. We can further compress the embedding and LM head layers, but leave this to future work due to inconsistencies in tokenizer vocabularies.
-Base Model Size $\Delta$ Size Comp. Factor
-Llama 2-7B 13.48 GB 1.24 GB 10.87
-Llama 2-13B 26.03 GB 2.09 GB 12.45
-Llama 2-70B 137.95 GB 8.95 GB 15.41
-Mistral-7B v0.1 14.48 GB 1.30 GB 11.14
 4.3 Latency Improvement
-To illustrate the idea of translating memory saving into improved latency, we implement a simple Triton kernel for GEMM with a binary matrix and scaling factor, as in BitDelta.
-We assume there are $B$ fine-tuned models served on the same GPU.
-For simplicity, we consider the setting where the requests are uniformly distributed, i.e., each model receives one request simultaneously.
+For simplicity, we consider the setting where each model receives one distinct request simultaneously.
+It would be insightful to develop more sophisticated serving systems, which we leave to future work.
 Following the decomposition in Eq.
-(6), the kernel is used to compute the batched matrix multiplication between $B$ binary matrices ($N\times M$) and $B$ high-precision activations ($L\times N$) where $N,M$ are intermediate dimensions and $L$ is the sequence length.
-We focus on decoding latency as opposed to the prefill latency, as it consumes most of the time.
-Tokens are generated one by one in decoding, meaning $L$ is always 1.
+(6), the $W_{INT1}A_{FP16}$ kernel is used to compute the batched matrix multiplication between $B$ binary matrices ($N\times M$) and $B$ high-precision activations ($L\times N$) where $N,M$ are intermediate dimensions and $L$ is the sequence length.
+We focus on decoding latency which dominates runtime, as opposed to prefill latency.
+Tokens are generated one by one when decoding, meaning $L$ is always 1.
+For all latency experiments we use a single A100 80GB with power limit set to 500W.
+Figure 4: Decoding latency of a linear layer, as in Eqn.
+6.
+Black: Shared base weight backbone $W_{\text{base}}X$.
+Blue: Batched activation-product with $B$ 1-bit deltas, as in BitDelta.
+Red: Batched activation-product with $B$ low-rank deltas, as in S-LoRA.
+Left: Ablation over hidden size, assuming $N=M$ and $B=1$.
+Right: Ablation over batch size, assuming $N=M=4096$.
 Kernel latency.
-We benchmark the decoding latency of our kernel, a batched linear operation over multiple deltas with a single base model, as in Eq.
-(6), corresponding to a single linear layer, and compare against naively computing the forward pass separately for each model.
-We ablate across the batch size and hidden size dimensions and find that our kernel consistently achieves about 2$\times$ speedup.
-Figure 4: Decoding latency of a linear layer with and without BitDelta.
-Blue: Naive forward pass with $B$ distinct fine-tuned models.
-Yellow: Batched forward pass with BitDelta, corresponding to one base model and $B$ 1-bit deltas, utilizing a Triton kernel.
-Left: Ablation over hidden size, assuming $N=M$ and $B=8$.
-Right: Ablation over batch size, assuming $N=M=8192$.
-End-to-end latency.
-We also benchmark the end-to-end decoding latency on Llama 2-7B variants with an input length of 128 (we find the decoding latency is less sensitive to the input length), ablated across the batch size.
-We find that BitDelta introduces overhead when the batch size is low.
-However, BitDelta scales better and successfully translates the saved GPU memory to improved decoding latency, starting at $B=4$.
-This is exacerbated at higher batch sizes, where the naive approach succumbs to out-of-memory issues and BitDelta is still performant.
-Figure 5: End-to-end decoding latency of Llama 2-7B variants with and without BitDelta.
+We benchmark the decoding latency of our kernel, a batched linear operation over multiple 1-bit deltas, corresponding to the delta component of Eq.
+(6).
+We compare this to the S-LoRA kernel, a batched linear operation over multiple low-rank deltas, and also compare this to the base weight backbone shared over all deltas.
+We set $r=128$ for S-LoRA, to maintain memory equivalence with BitDelta at $N=M=4096$.
+We profile the latency of the backbone ($W_{\text{base}}X$) and deltas ($\Delta X$) separately.
+Although $X$’s memory footprint scales with batch size, it is negligible compared to $W_{\text{base}}$, which remains constant.
+For typical low to medium batch settings, which is typical for $B\times N\ll N\times M$.
+In such settings, the overall memory footprint of the backbone is effectively independent of batch size, as shown in Figure 4 (left).
+This is in contrast with that of the deltas, which scales with the batch size, as each additional client in the batch adds an additional delta.
+At batch size 1 (Figure 4, right), backbone latency dominates over delta latency (BitDelta and S-LoRA) due to $W_{\text{base}}$’s 16$\times$ larger memory footprint compared to a single delta.
+As the batch size increases (Figure 4, left), the combined memory footprint of multiple deltas exceeds $W_{\text{base}}$ around $B=6$ to $B=8$.
+BitDelta underperforms slightly compared to S-LoRA in large-batch settings as the LoRA kernel is highly optimized for GPU.
+We emphasize that closing or even surpassing the gap is tractable.
+For example, Ma et al.
+[35] point out that $W_{INT1}A_{FP16}$ requires no multiplication operations and that new hardware can be co-designed with this in mind to drastically reduce energy/latency costs.
+Figure 5: Memory usage of Llama 2-7B, assuming each sequence in the batch has a length of $128$.
+Blue: Memory usage of the naive method, separately storing $B$ distinct fine-tuned models.
+Orange: Projected values for the naive method.
+Green: Memory usage of BitDelta.
+The naive forward pass succumbs to GPU memory issues at higher batch sizes.
+Figure 6: End-to-end decoding latency of Llama 2-7B.
 Blue: Naive forward pass with $B$ distinct fine-tuned models.
@@
 Green: Batched forward pass with BitDelta.
-The naive forward pass succumbs to GPU memory issues at higher batch sizes, whereas BitDelta is still performant.
-5 Conclusion and Discussion
-We propose BitDelta, a simple yet effective approach for efficiently quantizing the weight delta arising from fine-tuning in LLMs down to 1 bit.
+Gray: Batched forward pass with S-LoRA.
+The naive forward pass succumbs to GPU memory issues at higher batch sizes.
+End-to-end latency.
+We benchmark the end-to-end decoding latency on Llama 2-7B variants with an input length of 128 (we find the decoding latency is less sensitive to the input length), ablated across batch size.
+For BitDelta and S-LoRA, the forward pass consists of the addition of two components: a single backbone pass (batch independent) and a delta pass (scales with batch size).
+We compare BitDelta and S-LoRA with a naive method that computes each $W_{i}X_{i}$ separately in the forward pass.
+This naive approach scales poorly with batch size as it effectively maintains a separate backbone ($W_{i}$) for each client in the batch.
+Given the substantial memory footprint of the backbone, this leads to significant memory usage as batch size increases.
+In contrast, BitDelta and S-LoRA share a single backbone across all clients in the batch, with only the 16$\times$ smaller deltas scaling with batch size.
+This allows for more efficient memory utilization and better performance at larger batch sizes.
+We find that BitDelta and S-LoRA introduce overhead when the batch size is low.
+However, BitDelta and S-LoRA scale better and successfully translate the saved GPU memory to improved decoding latency, starting at $B=2$.
+This is exacerbated at larger batch sizes, where the naive approach succumbs to out-of-memory issues and BitDelta and S-LoRA are still performant.
+In the $B\geq 16$ regime, used in modern serving solutions, BitDelta has a >10×10$\times$10 × lower per-user decoding latency than the naive method.
+5 Conclusion
+We propose BitDelta, a simple but effective approach to efficiently quantifyings the weight delta arising from the fine-tuning of LLMs down to 1 bit.
 BitDelta encodes the sign bits of the weight delta and a per-weight matrix scaling factor, which is calibrated further through distillation.
@@
 BitDelta is fast and accurate, showcasing minimal performance degradation, and opens new avenues for efficient model deployment and resource utilization in machine learning.
-Future work.
-We mainly focus on the qualitative evaluation of BitDelta, as quality is the main challenge that most quantization methods struggle with.
-We show BitDelta can maintain the quality of the original fine-tuned models and further illustrate the idea of how to translate the memory saving into improved latency with a fused kernel.
-Yet, there is significant room for improvement – our kernel is fairly slow compared to the theoretical limit, considering the small memory footprint of binary matrices.
-It may also be possible to achieve better model quality by preserving salient weight deltas.
-Finally, the idea of calibrating certain scale factors through distillation may be applied more generally to PTQ methods.
-We hope our work motivates future refinement of scale factors, making low-bit quantized LLMs more robust.
-6 Impact Statement
-Environmental Sustainability and Cost Reduction.
-The reduction in GPU memory requirements through BitDelta translates to lower energy consumption and a reduction in costs associated with serving multiple fine-tuned models.
-By significantly lowering hardware requirements, BitDelta contributes to making AI technologies more eco-friendly and economically viable.
-This approach aligns with the growing need for sustainable and cost-effective computing solutions in the deployment of large-scale AI technologies.
+Acknowledgments and Disclosure of Funding
+We thank Together AI, MyShell AI, National Science Foundation (NSF), MIT-IBM Watson AI Lab, and MIT Amazon Science Hub for supporting this research.
+Appendix A Appendix
+A.1 Societal Impact
 Democratization of Fine-tuned Models.
@@
 We encourage further work on evaluation techniques to detect alignment loss in BitDelta, which can lead to the creation of robust methods for its mitigation.
-Acknowledgments
-We thank Together AI, MyShell AI, National Science Foundation (NSF), MIT-IBM Watson AI Lab, and MIT Amazon Science Hub for supporting this research.
-Appendix A Appendix
-Table 6: Full results of the comparison of BitDelta to the SVD based method, corresponding to Table 1.
-Model/Method Train Loss ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average normal-$\uparrow$ MT-Bench
+A.2 Additional Experiments
+Table 7: We train a $r=16$ LoRA finetune of Llama 2-7B on 1 epoch of UltraChat [17] and apply BitDelta with minimal performance degradation.
+This further shows the generality of BitDelta, which works on parameter-efficient fine-tunes in addition to full-parameter fine-tunes.
+Model/Method ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average $\uparrow$ MT-Bench
 Llama 2-7B 52.56 33.76 78.96 38.96 68.39 68.98 13.57 50.74
-Vicuna-7B v1.5 53.92 37.14 77.45 50.36 64.41 69.61 19.03 53.13 6.04
-BitDelta-Initial 0.41 54.69 36.74 78.47 47.63 66.31 68.75 19.56 53.16 5.67
-BitDelta 0.052 54.27 36.57 77.9 49.97 65.2 69.46 20.17 53.36 5.99
-SVD-Initial 0.78 54.61 33.96 78.61 44.56 67.77 68.03 15.92 51.92 4.45
-SVD 0.34 53.92 34.51 77.95 44.55 66.35 68.27 14.18 51.39 4.98
+Llama 2-7B UltraChat 54.52 34.14 78.99 46.84 70.83 69.53 14.71 52.79 4.93
+BitDelta 54.61 34.28 79.10 46.60 70.58 69.30 15.16 52.80 4.87
 Table 8: Full results of the application of BitDelta to quantized base models, corresponding to Table 6.
-Base Model Method ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average normal-$\uparrow$ MT-Bench
+Base Model Method ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average $\uparrow$ MT-Bench
 Baseline FP16 53.58 33.84 78.58 45.32 66.58 66.46 22.74 52.44 6.56
@@
 Table 9: Full results of the ablation over the fidelity of $\Delta$, corresponding to Figure 3.
-# bits in $\Delta$ ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average normal-$\uparrow$
+# bits in $\Delta$ ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average $\uparrow$
 Llama 2-7b 52.56 33.76 78.96 38.96 68.39 68.98 13.57 50.74
@@
 Table 10: Full results of BitDelta applied to fine-tuned models in the Llama-2 and Mistral families, corresponding to Table 2.
-Model Method ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average normal-$\uparrow$ MT-Bench $\uparrow$
+Model Method ARC BBH HellaSwag TruthfulQA LAMBADA WinoGrande GSM8K Average $\uparrow$ MT-Bench $\uparrow$
 Llama 2-7B 52.56 33.76 78.96 38.96 68.39 68.98 13.57 50.74